OpenStudy (anonymous):

the maturity value of a $16,000 non-interest-bearing ,simple discount 6 percent , 60 days not is ????

OpenStudy (gw2011):

(16,000)(0.06)(60/360)=(16,000)(0.06)(1/6)=$160. This is the discount amount.